OpenStudy (anonymous):

solve logarithmic function: ln(2x-3)=-2

OpenStudy (anonymous):

x=1.56767

OpenStudy (anonymous):

put "e" raised to both sides then you end up with 2x-3=-2^e and from there you use normal algebra to solve (but you will need a calc to figure out exact value)

OpenStudy (anonymous):

2x-3=e^-2 2x=e^2+3 x=(e^2+3)/2